# Neutrino flavor instabilities in neutron star mergers with moment transport: slow, fast, and collisional modes

### v1

@@ IV.1.1 Multi-energy results
The results of the multi-energy LSA with this isotropy assumption are shown in Fig. 2. We restrict our attention to the innermost zone where collisions can have a significant impact. The CFI growth rate is shown on the top panel, while the real part of $\Omega^{\prime}_{\mathrm{max}}$ is displayed on the bottom panel. There are clearly regions of gapless CFI (dark blue, $\mathrm{Re}(\Omega^{\prime}_{\mathrm{max}})\simeq 0$) and gapped CFI (yellow, $|\mathrm{Re}(\Omega^{\prime}_{\mathrm{max}})|\simeq|\Delta_{\mathcal{N}}-\overline{\Delta}_{\mathcal{N}}|\sim 10^{10}\ \mathrm{s}^{-1}$). The transitions between these regions are associated with larger growth rates (see the darker lines on the top panel), because they correspond to a “resonancelike” regime [41] where the number densities of $\nu_{e}$ and $\bar{\nu}_{e}$ are very close (see top left panel of Fig. 1). However, we will show in the following that these CFI regions are likely to harbor electron lepton number (ELN) crossings, such that a FFI will overwhelm even a resonancelike CFI. Overall, the CFI, estimated by this approximate isotropy assumption, appears to occur in large regions of the NSM snapshot, especially in the tidal arms, and just above and below the HMNS.
@@
Figure 3: Comparison of various determination of the isotropic and homogeneous CFI growth rate: multi-energy analysis (top left), single-energy analysis with averaged collision rates via method A (top right) and method B (bottom right). Two colormaps are used for the gapped (blue) and gapless (orange) modes. On the bottom left panel, the combination of monochromatic results (35) shows excellent agreement with the multi-energy results.
@@ IV.1.2 Comparison with energy-averaged approaches
This question was recently answered in the negative in Ref. [47] for the case of CCSNe, where it was shown that the energy averaging methods previously used (see Sec. II.5) largely overestimated the occurrence of CFI. While NSMs have comparable conditions to CCSNe, the differences are such that the same conclusion does not necessarily hold. To determine the answer for NSMs, we compare in Fig. 3 the different LSA approaches. The top left panel is the multi-energy result, identical to the top panel of Fig. 2 but where we use different color schemes for the gapped (blue) and gapless (orange) regions. We adopt the same color convention for the energy-averaged approaches: method A [see Eq. (19)] on the top right panel, and method B [see Eq. (21)] on the bottom right panel.
@@
We have shown in Eq. (18) that we expect the gapped regions to be correctly described by method A. This is indeed the case (compare the blue regions of the top panels of Fig. 3). However, method A predicts very large regions of gapless instability that are not present in the multi-energy analysis. We highlight in particular that the averaging method of Eq. (19) does not guarantee that the effective damping rates $\langle\Gamma\rangle_{\mathrm{A}},\langle\overline{\Gamma}\rangle_{\mathrm{A}}$ are positive. We illustrate this in Fig. 4, where we show the sign of $\langle\Gamma\rangle_{\mathrm{A}}+\langle\overline{\Gamma}\rangle_{\mathrm{A}}$ (we note that this expression is dominated by $\langle\Gamma\rangle_{\mathrm{A}}$). There are very large regions of negative energy-averaged damping rates, which coincide with the fictitious gapless modes visible on the top right panel of Fig. 3.
@@
The reader might be surprised by the fact that, if there are no $\nu_{x}$, methods A and B are identical—therefore, why wouldn’t the previous argument also apply to the gapless modes of method A? The reason is that “neglecting $\nu_{x}$” is not equivalent in both methods. In the derivation above, we neglect $\nu_{x}$ in several places. First, in $\Delta_{\mathcal{N}}$, which requires $g_{\nu_{e}}T_{\nu_{e}}^{3}\gg g_{\nu_{x}}T_{\nu_{x}}^{3}$. In the multi-energy instability criterion (33), since $\Gamma\propto p^{2}$ appears in the denominator, we need $g_{\nu_{e}}T_{\nu_{e}}\gg g_{\nu_{x}}T_{\nu_{x}}$. If we were to use method A, when computing $\langle\Gamma\rangle_{\mathrm{A}}$ with Eq. (19), since $\Gamma$ appears in the numerator, we would need $g_{\nu_{e}}T_{\nu_{e}}^{5}\gg g_{\nu_{x}}T_{\nu_{x}}^{5}$ before we are able to neglect the $\nu_{x}$ contribution. However, since $T_{\nu_{x}}>T_{\nu_{e}}$, but not by too much (see top left panel of Fig. 1), it is possible to neglect the $\nu_{x}$ contribution in Eqs. (31)–(33), but it is less straightforward in the expression of $\langle\Gamma\rangle_{\mathrm{A}}$, and this is what leads to the negative sign encountered in Fig. 4.
@@
To assess more quantitatively the performance of the energy-averaged approaches, we show in Fig. 5 the cumulative distribution of the difference between the monochromatic estimate and the multi-energy result, $\delta[\mathrm{Im}(\Omega)_{\mathrm{max}}]=|\mathrm{Im}(\Omega)_{\text{avE}}/\mathrm{Im}(\Omega)_{\text{multiE}}-1|$, distinguishing between gapped and gapless regions. For the gapped regions, we compare the multi-energy growth rate with method A, proving that the relative difference is smaller than a percent for $80\%$ of the grid points, as expected. For the gapless regions we use the method B monochromatic result. The performance is clearly worse than in the gapped case, however, the relative difference is typically of order unity. This is confirmed by looking more specifically at the ratio between the method B prediction and the actual multi-energy gapless CFI growth rate. The histogram of this ratio is shown in Fig. 6. Although there is some dispersion, we see that for the conditions encountered in this NSM snapshot, there is typically a factor 2 overestimation of the growth rate by method B. This is very different from what was observed in CCSN environments in [47], and we attribute it to the hierarchy of number densities $\mathcal{N}_{xx}=\overline{\mathcal{N}}_{xx}\ll\mathcal{N}_{ee},\overline{\mathcal{N}}_{ee}$ (see Appendix C).
Finally, we note that in the monochromatic case away from the resonance regime, gapped and gapless instabilities are mutually exclusive, as can be seen by comparing Eqs. (49) and (50): assuming for instance $\Delta_{\mathcal{N}}>\overline{\Delta}_{\mathcal{N}}$, $\mathrm{Im}(\Omega^{\prime}_{-})>0$ if $\Gamma_{\mathcal{N}}/\overline{\Gamma}_{\mathcal{N}}<\overline{\Delta}_{\mathcal{N}}/\Delta_{\mathcal{N}}<1$, while $\mathrm{Im}(\Omega^{\prime}_{+})>0$ if $\Gamma_{\mathcal{N}}/\overline{\Gamma}_{\mathcal{N}}>\Delta_{\mathcal{N}}/\overline{\Delta}_{\mathcal{N}}>1$. However, there is no such mutual exclusion in the multi-energy case, and we have found instances where both a gapless and a gapped mode are unstable. This can be identified when simultaneously the gapped mode of method A and the gapless mode of method B are unstable (i.e., a point in Fig. 3 in blue in the top right panel and in orange in the bottom right panel). In this situation, the multi-energy result shown on the top left panel corresponds to the largest growth rate among the two. Since we know that the gapped method A growth rate matches the multi-energy gapped one, and the gapless method B growth rate must be typically divided by two, we build the following monochromatic estimate of the multi-energy CFI:
@@
This combination is depicted on the bottom left panel of Fig. 3, and shows excellent qualitative and quantitative agreement with the multi-energy results.
@@ IV.2 Many-body corrections to the absorption rates
These corrections are jointly determined by the kinematics of neutrino-nucleon reactions, the thermal distribution, the dispersion relation and the residual interactions of nucleons, which all vary with the nucleon density, local temperature, proton fraction and incoming neutrino energies. At low densities ($\lesssim 10^{10}\,\mathrm{g\,cm^{-3}}$), the many-body corrections are not sensitive to the nucleon interactions and are mainly determined by the kinematics. Correctly applying the kinematics automatically includes the phase space corrections, which depend on the incoming neutrino energy and the temperature, and are slightly different for $\nu_{e}$ and $\bar{\nu}_{e}$. At medium ($10^{10}\,\mathrm{g\,cm^{-3}}\lesssim\rho\lesssim 10^{13}\,\mathrm{g\,cm^{-3}}$) and high ($\rho\gtrsim 10^{13}\,\mathrm{g\,cm^{-3}}$) densities, the nucleon-nucleon interactions noticeably alter the correlations of the nucleonic matter and change the neutrino opacity. Here, we apply the Hartree-Fock + Random Phase Approximation (HF+RPA) as in [110] to describe the correlation effects. The mean-field (MF) effect described by HF is mainly caused by the difference of neutron and proton potentials, which increases (decreases) the opacity of $\nu_{e}$ ($\bar{\nu}_{e}$). On the other hand, the RPA effect significantly modifies the differential neutrino-nucleon cross sections and may result in a resonance peak. Recent studies [111, 38] show that in NSM- and CCSN-like conditions, the RPA effect may decrease (increase) the opacity of $\nu_{e}$ ($\bar{\nu}_{e}$), countering the MF effect. At medium densities, the MF and the RPA effect are approximately comparable, resulting in a relatively weak net effect. At high densities, the MF effect may dominate the RPA effect at low temperature and low proton fraction, changing the opacities by several orders of magnitude. When the proton fraction is approaching $50\,\%$, the increase (decrease) of $\kappa_{\mathrm{a},e}$ ($\bar{\kappa}_{\mathrm{a},e}$) due to the MF effect is significantly weakened. At high temperature (e.g. $T\gtrsim 20\,\mathrm{MeV}$), the many-body correction is much less important than at low temperature. The standard opacities based on Bruenn et al. [112] include a simplified description of many-body corrections which may overestimate the reduction of $\kappa_{\mathrm{a},e},\,\bar{\kappa}_{\mathrm{a},e}$ at high temperatures, such that in that regime our modified opacities are larger than the standard ones.
@@
Figure 7: Effect of many-nucleon corrections on the absorption opacities and the homogeneous and isotropic CFI in the 7 ms postmerger snapshot. Top: ratio of modified (“mod”) and original (“orig”) energy-averaged opacities [Eq. (20)] for $\nu_{e}$. Middle: same for $\bar{\nu}_{e}$. Bottom: ratio of CFI growth rates with modified and original opacities across the snapshot. The original CFI growth rates are shown on the top panel of Fig. 2. The turquoise color represents new regions of instability originating from the modified opacities, and the pink color represents stabilization of previously unstable regions. Note that we used the dispersion relation approach discussed in Appendix B to find the instabilities. It agrees perfectly with the stability matrix results (see also the comparison in Sec. B.3).
@@
These effects are at the origin of the changes in opacities depicted on the first two panels of Fig. 7. In order to represent in a single quantity the net effect of the many-body corrections, we plot the ratio of the gray opacities as defined in Eq. (20). Note that at densities above the saturation density, we do not calculate the many-body correction (white region in the HMNS), since neutrinos are in classical equilibrium, for which there is no instability [77]. For electron neutrinos (top panel), we notice a general trend of decrease in $\langle\kappa_{\mathrm{a},e}\rangle$, except for a sharp increase close to the HMNS—but this happens in a region that is far from collisional instability. We also notice sharp changes in the radial direction, which correspond to sharp changes in the thermodynamic quantities, associated with the tidal arms around the HMNS. For electron antineutrinos, we get a similar pattern but with generally increasing opacities, and a sharp change in the ratio around the same region as for the electron neutrino opacities. Near the remnant and when the density is still small enough that we calculate the correction, the increase in $\langle\bar{\kappa}_{\mathrm{a},e}\rangle$ is consistent with the mechanisms outlined above in the high temperature regime.
@@
We check how these corrections modify the collisional instability landscape by plotting on the bottom panel of Fig. 7 the ratio of the CFI growth rate with and without many-body corrections (compare with Fig. 2). We can see that for the 7 ms snapshot we study, these corrections do not significantly alter the unstable regions. Some very low unstable growth rates, near the fringes of instability, seem to vanish (pink color), and some sparse new unstable regions appear (turquoise color). The ratios of the modified to original growth rates for the rest of the unstable regions are mostly <1<1, but within order of unity. Since, in the snapshot, neutrino opacities are mostly larger than the antineutrino opacities (see bottom right panel of Fig. 1), the decrease (increase) in $\langle{\kappa}_{\mathrm{a},e}\rangle$ ($\langle\bar{\kappa}_{\mathrm{a},e}\rangle$), reduces the discrepancy between the collision rates, and thus subdues the instability. These relatively small changes in growth rates are consistent with the variations of opacities that we can see in the CFI regions, which are typically below $\sim 20\,\%$ (see top and middle panels of Fig. 7). Since there are no dramatic changes due to the many-body corrections, for consistency with the collision rates that were used in the NSM simulation from which we extract the data, we will use the original opacities in the rest of this work.

### v2

@@ IV.1.1 Multi-energy results
The results of the multi-energy LSA with this isotropy assumption are shown in Fig. 2. We restrict our attention to the innermost zone where collisions can have a significant impact. The CFI growth rate is shown in the top panel, while the real part of $\Omega^{\prime}_{\mathrm{max}}$ is displayed in the bottom panel. There are clearly regions of gapless CFI (dark blue, $\mathrm{Re}(\Omega^{\prime}_{\mathrm{max}})\simeq 0$) and gapped CFI (yellow, $|\mathrm{Re}(\Omega^{\prime}_{\mathrm{max}})|\simeq|\Delta_{\mathcal{N}}-\overline{\Delta}_{\mathcal{N}}|\sim 10^{10}\ \mathrm{s}^{-1}$). The transitions between these regions are associated with larger growth rates (see the darker lines on the top panel), because they correspond to a “resonancelike” regime [41] where the number densities of $\nu_{e}$ and $\bar{\nu}_{e}$ are very close (see top left panel of Fig. 1). However, we will show in the following that these CFI regions are likely to harbor electron lepton number (ELN) crossings, such that a FFI will overwhelm even a resonancelike CFI. Overall, the CFI, estimated by this approximate isotropy assumption, appears to occur in large regions of the NSM snapshot, especially in the tidal arms, and just above and below the HMNS.
@@
Figure 3: Comparison of various determination of the isotropic and homogeneous CFI growth rate: multi-energy analysis (top left), single-energy analysis with averaged collision rates via method A (top right) and method B (bottom right). Two color maps are used for the gapped (blue) and gapless (orange) modes. In the bottom left panel, the combination of monochromatic results (35) shows excellent agreement with the multi-energy results.
@@ IV.1.2 Comparison with energy-averaged approaches
This question was recently answered in the negative in Ref. [47] for the case of CCSNe, where it was shown that the energy-averaging methods previously used (see Sec. II.5) largely overestimated the occurrence of CFI. While NSMs have comparable conditions to CCSNe, the differences are such that the same conclusion does not necessarily hold. To determine the answer for NSMs, we compare in Fig. 3 the different LSA approaches. The top left panel is the multi-energy result, identical to the top panel of Fig. 2 but where we use different color schemes for the gapped (blue) and gapless (orange) regions. We adopt the same color convention for the energy-averaged approaches: method A [see Eq. (19)] in the top right panel and method B [see Eq. (21)] in the bottom right panel.
@@
We have shown in Eq. (18) that we expect the gapped regions to be correctly described by method A. This is indeed the case (compare the blue regions of the top panels of Fig. 3). However, method A predicts very large regions of gapless instability that are not present in the multi-energy analysis. We highlight in particular that the averaging method of Eq. (19) does not guarantee that the effective damping rates $\langle\Gamma\rangle_{\mathrm{A}}$ and $\langle\overline{\Gamma}\rangle_{\mathrm{A}}$ are positive. We illustrate this in Fig. 4, where we show the sign of $\langle\Gamma\rangle_{\mathrm{A}}+\langle\overline{\Gamma}\rangle_{\mathrm{A}}$ (we note that this expression is dominated by $\langle\Gamma\rangle_{\mathrm{A}}$). There are very large regions of negative energy-averaged damping rates, which coincide with the fictitious gapless modes visible in the top right panel of Fig. 3.
@@
The reader might be surprised by the fact that, if there are no $\nu_{x}$, methods A and B are identical—therefore, why would the previous argument not also apply to the gapless modes of method A? The reason is that “neglecting $\nu_{x}$” is not equivalent in both methods. In the derivation above, we neglect $\nu_{x}$ in several places. First, in $\Delta_{\mathcal{N}}$, which requires $g_{\nu_{e}}T_{\nu_{e}}^{3}\gg g_{\nu_{x}}T_{\nu_{x}}^{3}$. In the multi-energy instability criterion (33), since $\Gamma\propto p^{2}$ appears in the denominator, we need $g_{\nu_{e}}T_{\nu_{e}}\gg g_{\nu_{x}}T_{\nu_{x}}$. If we were to use method A, when computing $\langle\Gamma\rangle_{\mathrm{A}}$ with Eq. (19), since $\Gamma$ appears in the numerator, we would need $g_{\nu_{e}}T_{\nu_{e}}^{5}\gg g_{\nu_{x}}T_{\nu_{x}}^{5}$ before we are able to neglect the $\nu_{x}$ contribution. However, since $T_{\nu_{x}}>T_{\nu_{e}}$, but not by too much (see top left panel of Fig. 1), it is possible to neglect the $\nu_{x}$ contribution in Eqs. (31)–(33), but it is less straightforward in the expression of $\langle\Gamma\rangle_{\mathrm{A}}$, and this is what leads to the negative sign encountered in Fig. 4.
@@
To assess more quantitatively the performance of the energy-averaged approaches, we show in Fig. 5 the cumulative distribution of the difference between the monochromatic estimate and the multi-energy result, $\delta[\mathrm{Im}(\Omega)_{\mathrm{max}}]=|\mathrm{Im}(\Omega)_{\text{avE}}/\mathrm{Im}(\Omega)_{\text{multiE}}-1|$, distinguishing between gapped and gapless regions. For the gapped regions, we compare the multi-energy growth rate with method A, proving that the relative difference is smaller than a percent for $80\%$ of the grid points, as expected. For the gapless regions we use the method B monochromatic result. The performance is clearly worse than in the gapped case; however, the relative difference is typically of order unity. This is confirmed by looking more specifically at the ratio between the method B prediction and the actual multi-energy gapless CFI growth rate. The histogram of this ratio is shown in Fig. 6. Although there is some dispersion, we see that for the conditions encountered in this NSM snapshot, there is typically a factor 2 overestimation of the growth rate by method B. This is very different from what was observed in CCSN environments in [47], and we attribute it to the hierarchy of number densities $\mathcal{N}_{xx}=\overline{\mathcal{N}}_{xx}\ll\mathcal{N}_{ee},\overline{\mathcal{N}}_{ee}$ (see Appendix C).
Finally, we note that in the monochromatic case away from the resonance regime, gapped and gapless instabilities are mutually exclusive, as can be seen by comparing Eqs. (49) and (50): assuming for instance $\Delta_{\mathcal{N}}>\overline{\Delta}_{\mathcal{N}}$, $\mathrm{Im}(\Omega^{\prime}_{-})>0$ if $\Gamma_{\mathcal{N}}/\overline{\Gamma}_{\mathcal{N}}<\overline{\Delta}_{\mathcal{N}}/\Delta_{\mathcal{N}}<1$, while $\mathrm{Im}(\Omega^{\prime}_{+})>0$ if $\Gamma_{\mathcal{N}}/\overline{\Gamma}_{\mathcal{N}}>\Delta_{\mathcal{N}}/\overline{\Delta}_{\mathcal{N}}>1$. However, there is no such mutual exclusion in the multi-energy case, and we have found instances where both a gapless and a gapped mode are unstable. This can be identified when simultaneously the gapped mode of method A and the gapless mode of method B are unstable (i.e., a point in Fig. 3 in blue in the top right panel and in orange in the bottom right panel). In this situation, the multi-energy result shown in the top left panel corresponds to the largest growth rate among the two. Since we know that the gapped method A growth rate matches the multi-energy gapped one, and the gapless method B growth rate must be typically divided by 2, we build the following monochromatic estimate of the multi-energy CFI:
@@
This combination is depicted in the bottom left panel of Fig. 3, and shows excellent qualitative and quantitative agreement with the multi-energy results.
@@ IV.2 Many-body corrections to the absorption rates
These corrections are jointly determined by the kinematics of neutrino-nucleon reactions, the thermal distribution, the dispersion relation and the residual interactions of nucleons, which all vary with the nucleon density, local temperature, proton fraction and incoming neutrino energies. At low densities ($\lesssim 10^{10}\,\mathrm{g\,cm^{-3}}$), the many-body corrections are not sensitive to the nucleon interactions and are mainly determined by the kinematics. Correctly applying the kinematics automatically includes the phase space corrections, which depend on the incoming neutrino energy and the temperature, and are slightly different for $\nu_{e}$ and $\bar{\nu}_{e}$. At medium ($10^{10}\,\mathrm{g\,cm^{-3}}\lesssim\rho\lesssim 10^{13}\,\mathrm{g\,cm^{-3}}$) and high ($\rho\gtrsim 10^{13}\,\mathrm{g\,cm^{-3}}$) densities, the nucleon-nucleon interactions noticeably alter the correlations of the nucleonic matter and change the neutrino opacity. Here, we apply the Hartree-Fock + random phase approximation (HF+RPA) as in [111] to describe the correlation effects. The mean-field (MF) effect described by HF is mainly caused by the difference of neutron and proton potentials, which increases (decreases) the opacity of $\nu_{e}$ ($\bar{\nu}_{e}$). On the other hand, the RPA effect significantly modifies the differential neutrino-nucleon cross sections and may result in a resonance peak. Recent studies [112, 38] show that in NSM- and CCSN-like conditions, the RPA effect may decrease (increase) the opacity of $\nu_{e}$ ($\bar{\nu}_{e}$), countering the MF effect. At medium densities, the MF and the RPA effect are approximately comparable, resulting in a relatively weak net effect. At high densities, the MF effect may dominate the RPA effect at low temperature and low proton fraction, changing the opacities by several orders of magnitude. When the proton fraction is approaching $50\,\%$, the increase (decrease) of $\kappa_{\mathrm{a},e}$ ($\bar{\kappa}_{\mathrm{a},e}$) due to the MF effect is significantly weakened. At high temperature (e.g., $T\gtrsim 20\,\mathrm{MeV}$), the many-body correction is much less important than at low temperature. The standard opacities based on Bruenn [113] include a simplified description of many-body corrections which may overestimate the reduction of $\kappa_{\mathrm{a},e},\,\bar{\kappa}_{\mathrm{a},e}$ at high temperatures, such that in that regime our modified opacities are larger than the standard ones.
@@
Figure 7: Effect of many-nucleon corrections on the absorption opacities and the homogeneous and isotropic CFI in the 7 ms postmerger snapshot. Top: ratio of modified (“mod”) and original (“orig”) energy-averaged opacities [Eq. (20)] for $\nu_{e}$. Middle: same for $\bar{\nu}_{e}$. Bottom: ratio of CFI growth rates with modified and original opacities across the snapshot. The original CFI growth rates are shown in the top panel of Fig. 2. The turquoise color represents new regions of instability originating from the modified opacities, and the pink color represents stabilization of previously unstable regions. Note that we used the dispersion relation approach discussed in Appendix B to find the instabilities. It agrees perfectly with the stability matrix results (see also the comparison in Sec. B.3).
@@
These effects are at the origin of the changes in opacities depicted in the first two panels of Fig. 7. In order to represent in a single quantity the net effect of the many-body corrections, we plot the ratio of the gray opacities as defined in Eq. (20). Note that at densities above the saturation density, we do not calculate the many-body correction (white region in the HMNS), since neutrinos are in classical equilibrium, for which there is no instability [78]. For electron neutrinos (top panel), we notice a general trend of decrease in $\langle\kappa_{\mathrm{a},e}\rangle$, except for a sharp increase close to the HMNS—but this happens in a region that is far from collisional instability. We also notice sharp changes in the radial direction, which correspond to sharp changes in the thermodynamic quantities, associated with the tidal arms around the HMNS. For electron antineutrinos, we get a similar pattern but with generally increasing opacities, and a sharp change in the ratio around the same region as for the electron neutrino opacities. Near the remnant and when the density is still small enough that we calculate the correction, the increase in $\langle\bar{\kappa}_{\mathrm{a},e}\rangle$ is consistent with the mechanisms outlined above in the high-temperature regime.
@@
We check how these corrections modify the collisional instability landscape by plotting in the bottom panel of Fig. 7 the ratio of the CFI growth rate with and without many-body corrections (compare with Fig. 2). We can see that for the 7 ms snapshot we study, these corrections do not significantly alter the unstable regions. Some very low unstable growth rates, near the fringes of instability, seem to vanish (pink color), and some sparse new unstable regions appear (turquoise color). The ratios of the modified to original growth rates for the rest of the unstable regions are mostly <1<1, but within order of unity. Since, in the snapshot, neutrino opacities are mostly larger than the antineutrino opacities (see bottom right panel of Fig. 1), the decrease (increase) in $\langle{\kappa}_{\mathrm{a},e}\rangle$ ($\langle\bar{\kappa}_{\mathrm{a},e}\rangle$), reduces the discrepancy between the collision rates, and thus subdues the instability. These relatively small changes in growth rates are consistent with the variations of opacities that we can see in the CFI regions, which are typically below $\sim 20\,\%$ (see top and middle panels of Fig. 7). Since there are no dramatic changes due to the many-body corrections, for consistency with the collision rates that were used in the NSM simulation from which we extract the data, we will use the original opacities in the rest of this work.